OpenStudy (anonymous):

write an equation in point-slope form for the line that contains the given points: (1,-6) and (2.-4) write an equation in point-slope form for the line that contains the given points: (1,-6) and (2.-4) @Mathematics

OpenStudy (nottim):

I might not answer, but please show others what oyu have done so far.

OpenStudy (anonymous):

i don't know how to do this, sorry

OpenStudy (anonymous):

find the slope first \[m=(y _{1}-y _{2})/(x _{1}-x _{2})\] you will get (-6-(-4))/(1-2) which turns out to be -2/-1 or just simply 2 use the point slope formula to get the equation. \[y-y _{1}=m(x-x _{1})\] you will get y+6=2(x-1)

OpenStudy (anonymous):

put it in y=mx+b and you will have your answer!

OpenStudy (anonymous):

is it -6?

OpenStudy (anonymous):

they are just asking you for the equation. not an answer. the equation is y=2x-8

OpenStudy (anonymous):

ohhhh ic

OpenStudy (anonymous):

can i ask you another question, familiar to the one i just asked?

OpenStudy (anonymous):

yes what is it?

OpenStudy (anonymous):

gimme a medal

OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that has the given slope and contains the given point: m= 2/3 (9,5)

OpenStudy (anonymous):

well do you know how to use the point slope formula? I posted it above, \[(y-y _{1})=m(x-x _{1})\] m=slope \[y _{1}\]\[= y coordinate\] \[x _{1}= x coordinate.\] if you plug in the correct values and do a little algebra you will get the equation of the line. let me know what you get.

OpenStudy (anonymous):

yes its y=mx+b

OpenStudy (anonymous):

5=2/3(9)+b is that correct so far?

OpenStudy (anonymous):

no. you have to put it into the formula that i posted above. NOT y=mx+b

OpenStudy (anonymous):

well now that i think of it you could do it that way. and solve for b but i havent done that in years.. and it seems like it is more complicated

OpenStudy (anonymous):

what did you get for an answer? i'll see if i can get the same answer using y=mx+b

OpenStudy (anonymous):

y=(2/3)x-1

OpenStudy (anonymous):

yea i can get the same thing using my formula